OpenStudy (anonymous):

find the differential of the dependent variable: g(x)= (x^2 + 3x +1)((1/x) + (1/x^2) +3)

OpenStudy (mathmale):

Fortunately, the "differential" is a relatively easy concept. If y=g(x), then (dy/dx) = g '(x), and this can be re-written as dy = g '(x) dx. So, in your math problem, to find the differential of the function g(x), write dy = g '(x)*dx, after you've found the derivative of g '(x).

OpenStudy (anonymous):

heeeeyy its you again! :D

OpenStudy (mathmale):

If y = x^2, the derivative is dy/dx = 2x, and the differential is dy = 2xdx. Re: "heeeeyy its you again!" Is that good? or...not? :)

OpenStudy (anonymous):

good ofc

OpenStudy (mathmale):

What a relief! :)

OpenStudy (anonymous):

so i do the derivative of each parenthesis?

OpenStudy (mathmale):

You're given the function g(x). Just find the derivative of g(x) in the usual way. Notice that your g(x) is a PRODUCT, so use the product rule.

OpenStudy (anonymous):

im not sure what the derivative of the fractions are

OpenStudy (mathmale):

The second multiplicand is \[((1/x) + (1/x^2) +3)\]

OpenStudy (mathmale):

...and it's probably best to rewrite this before taking the derivative:\[x ^{-1}+x ^{-2}+3\]

OpenStudy (mathmale):

From this point on, you just apply the power rule to each term. For example, the deriv. of the first term is -x^(-2). What are the derivs. of x^(-2) and 3?

OpenStudy (anonymous):

im sorry where did you get the -x^(-2) from?

OpenStudy (mathmale):

\[\frac{ d }{ dx }(\frac{ 1 }{ x })=\frac{ d }{ dx }x ^{-1}=-1x ^{-2}\]by the Power Rule.

OpenStudy (anonymous):

ah yes i forgot

OpenStudy (anonymous):

so itd be (2x+3) + (-x^-2 - x^-3)

OpenStudy (mathmale):

So, \[\frac{ d }{ dx }(\frac{ 1 }{ x }+\frac{ 1 }{ x^2 }+3) = -1x ^{-2} -2x ^{-3}+0\]

OpenStudy (anonymous):

ops forgot the 2

OpenStudy (mathmale):

Rememb\[\frac{ d }{ dx }u*v=u \frac{ dv }{ dx }+v \frac{ du }{ dx }\]er that the Product will crease a result with two terms and four factors:

OpenStudy (mathmale):

create (not crease)

OpenStudy (mathmale):

so that the derivative of your function g(x) will come out to (x^2+3x+1)(-x^(-2) -2x^(-3)) + ( )( ), two terms and four factors.

OpenStudy (mathmale):

\[\frac{ d }{ dx }=(x^2+3x+1)(-x ^{-2}-2x ^{-3}) + (~)*(~)\]

OpenStudy (anonymous):

yup got that

OpenStudy (mathmale):

can you fill in the two empty sets of parentheses? Good. so, supposing you end up with g '(x). Your differential will then be\[dy=dg=g '(x)*dx.\]

OpenStudy (anonymous):

so its dg(x) = (x^2 + 3x +1)(-x^-2 - 2x^-3) + (1/x + 1/x^2 +3) (2x+3) dx

OpenStudy (mathmale):

\[dg = [(x^2 + 3x +1)(-x^-2 - 2x^-3) + (1/x + 1/x^2 +3) (2x+3)] dx\]

OpenStudy (mathmale):

the only significant change I've made here was to insert brackets: [ ].

OpenStudy (anonymous):

thanks! how long will you be on OS for because i have quite a few questions and i was wondering if you may help me with them

OpenStudy (mathmale):

Not sure, really. I've spent a significant part of my day on OS, and realize I'm overdoing it. Nevertheless: go ahead and post one of your more challenging remaining questions.